# Gauge preservation in renormalization for Yang-Mills and gravity theories

### v2

@@ 3.3 The propagator in certain higher order gravity theories
$S_{g}=\frac{1}{16\pi G_{N}}\int\bigg{[}R+RF_{1}(\square)R+R_{\mu\nu}F_{2}(\square)R^{\mu\nu}+R_{\mu\nu\alpha\beta}F_{3}(\square)R_{\mu\nu\alpha\beta}\bigg{]}\sqrt{-g}d^{4}x+S_{gf}.$ (3.106)
@@
$P^{2}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\theta_{\alpha\nu}+\theta_{\beta\nu}\theta_{\alpha\mu})-\frac{1}{3}\theta_{\beta\alpha}\theta_{\mu\nu}$
$P^{1}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\omega_{\alpha\nu}+\theta_{\beta\nu}\omega_{\alpha\mu}+\theta_{\alpha\mu}\omega_{\beta\nu}+\theta_{\alpha\nu}\omega_{\beta\mu})$
@@
$=-k^{2}a(-k^{2})h^{\alpha\beta}(P^{2}_{\alpha\beta,\mu\nu}+P^{0}_{\alpha\beta,\mu\nu}+P^{1}_{\alpha\beta,\mu\nu}+P^{0-w}_{\alpha\beta,\mu\nu})h^{\mu\nu}$
@@
$D^{\alpha\beta,\mu\nu}=-\frac{1}{k^{2}}\bigg{[}\frac{2}{a}P^{2}+2P^{1}+\frac{2a+d-1}{2D}P^{0-s}+\frac{a+3d}{2D}P^{0-w}+\frac{\sqrt{3}}{4D}(P^{0-sw}+P^{0-ws})\bigg{]}_{\alpha\beta,\mu\nu}.$ (3.112)
@@ 3.4 Gravitational BRST symmetries
As for Yang-Mills fields, the gauge fixed action $S_{gf}$ is not invariant under the infinitesimal diffeomorphisms (3.100). In fact, it should not be so, as it reflects a particular choice of gauge. The task is to find a transformation, if it exists, leaving the gauge fixed action unchanged, perhaps up to a boundary term. This is the analog to the BRST symmetry described in (2.19)-(2.30), and one may try to generalize the procedure described in those formulas adapted to the present case. Instead of going through all the calculations as for the Yang-Mills case, one may employ the gained knowledge in those sections. A try may be to find a BRST transformation that increases the ghost number by one, supposing that these numbers are the same as for the Yang-Mills counterpart. In fact they are, as the assignation of the same ghost numbers than in Yang-Mills theories leads to a zero ghost number to the action $S_{gf}$, which is a desired feature. The ansatz for BRST transformation is then
$\delta h^{\mu\nu}=\kappa D_{\alpha}^{\mu\nu}\eta^{\alpha}\epsilon,\qquad\delta\eta^{\alpha}=-\kappa^{2}\partial_{\beta}\eta^{\alpha}\eta^{\beta}\epsilon,\qquad\delta\overline{\eta}_{\tau}=-\frac{\kappa^{2}}{\lambda}\square^{3}F_{\tau}\epsilon.$ (3.115)
@@
$\delta s\eta^{\alpha}=k^{4}\partial_{\beta}(\partial_{\gamma}\eta^{\alpha}\eta^{\gamma}\epsilon)\eta^{\beta}+\kappa^{4}\partial_{\beta}\eta^{\alpha}\partial_{\eta}\eta^{\beta}\eta^{\gamma}\epsilon$
@@
$-\frac{\kappa^{3}}{\lambda}(\square^{3}F_{\tau}\epsilon)F_{\mu\nu}^{\tau}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\overline{\eta}_{\tau}F_{\mu\nu}^{\tau}\delta(D^{\mu\nu}_{\alpha}\eta^{\alpha})$ (3.116)
@@
$\delta(D^{\mu\nu}_{\alpha}\eta^{\alpha})=D^{\mu\nu}_{\alpha}\partial_{\gamma}\eta^{\alpha}\eta^{\gamma}\epsilon-\frac{\delta D_{\alpha}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\gamma}^{\alpha\beta}\eta^{\alpha}\eta^{\gamma}\epsilon$
$=D^{\mu\nu}_{\alpha}\partial_{\gamma}\eta^{\alpha}\eta^{\gamma}\epsilon-\frac{1}{2}\bigg{(}\frac{\delta D_{\alpha}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\gamma}^{\alpha\beta}-\frac{\delta D_{\gamma}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\alpha}^{\alpha\beta}\bigg{)}\eta^{\alpha}\eta^{\gamma}\epsilon,$ (3.117)
@@
$(\delta_{\eta}\delta_{\xi}-\delta_{\xi}\delta_{\eta})h^{\mu\nu}=\bigg{(}\frac{\delta D_{\alpha}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\gamma}^{\alpha\beta}-\frac{\delta D_{\gamma}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\alpha}^{\alpha\beta}\bigg{)}\eta^{\alpha}\xi^{\gamma}.$
@@
$\bigg{(}\frac{\delta D_{\alpha}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\gamma}^{\alpha\beta}-\frac{\delta D_{\gamma}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\alpha}^{\alpha\beta}\bigg{)}\eta^{\alpha}\eta^{\gamma}=2D^{\mu\nu}_{\alpha}\partial_{\gamma}\eta^{\alpha}\eta^{\gamma}.$
@@ 3.5 The effective gravitational action
In analogy with the procedure in (2.33), the effective action $G(T,\xi,\overline{\xi},K,L)$ for gravitational theories, generating the connected correlation functions, is given by
@@
$\exp\{i\int d^{4}x[L_{s}-\frac{\kappa^{2}}{2\lambda}F_{\tau}\square F^{\tau}+\overline{\eta}_{\tau}F_{\mu\nu}^{\tau}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa K_{\mu\nu}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa^{2}L_{\alpha}\partial_{\beta}\eta^{\alpha}\eta^{\beta}]\}.$
Here the current $J_{\mu\nu}h^{\mu\nu}$ is the analogous to the gauge field current coupling $J_{\mu}A^{\mu}$ for Yang-Mills theories, and $J_{\mu\nu}$ may be identified as the energy momentum tensor $J_{\mu\nu}=T_{\mu\nu}$ for an external source. As in those theories, the sources $K_{\mu\nu}$ and $L_{\alpha}$ are coupled to the nilpotent part of the BRST invariance. From this expression it follows that the gauge fixed action for the gravitational perturbation is given by
@@ 3.6 Consequences of gravitational Slavnov-Taylor identiies
$\frac{\delta S_{(0)gf}}{\delta K_{\mu\nu}}=\kappa D^{\mu\nu}_{\alpha}\eta_{0}^{\eta},\qquad\frac{\delta S_{(0)gf}}{\delta h^{\mu\nu}}=(\kappa K_{\alpha\beta}-\overline{\eta}^{\tau}\overline{F}_{\alpha\beta}^{\tau})\frac{\delta D_{\gamma}^{\alpha\beta}}{\delta h_{0}^{\mu\nu}}\eta_{0}^{\gamma}+\frac{\delta S_{0}}{\delta h_{0}^{\mu\nu}},$
@@
Here the reversed arrow $\overleftarrow{}$ indicates the action of the operator on the left. With the help of the last two formulas, the solution $\Gamma_{1}=S+\int d^{4}x\sigma_{0}X$, can be found by use of (3.129) and (3.127), leading to
@@ 3.7 Physical interpretation of the counter terms for gravitational theories
$+\int d^{4}x\bigg{[}-(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})\frac{\partial P^{\rho\sigma}}{\delta h^{\mu\nu}}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa L_{\sigma}\bigg{(}\frac{\delta Q^{\sigma}_{\tau}}{\delta h^{\mu\nu}}\eta^{\tau}D_{\beta}^{\mu\nu}-\kappa Q^{\sigma}_{\tau}\partial_{\beta}(\eta^{\tau})\bigg{)}\eta^{\beta}\bigg{]}.$ (3.133)
@@
$\widetilde{\delta}(\widetilde{h}_{\mu\nu})-\delta(\widetilde{h}_{\mu\nu})=\widetilde{D}^{\mu\nu}_{\alpha}\widetilde{\eta}^{\alpha}\epsilon-D^{\mu\nu}_{\alpha}\eta^{\alpha}\epsilon-\frac{\partial P^{\rho\sigma}}{\delta h^{\gamma\delta}}D^{\gamma\delta}_{\alpha}\eta^{\alpha}\epsilon$
$=\bigg{(}\frac{\delta D^{\rho\sigma}_{\alpha}}{\delta h^{\mu\nu}}\eta^{\alpha}P^{\mu\nu}-D^{\rho\sigma}_{\alpha}(Q^{\alpha}_{\tau}\eta^{\tau})-\frac{\partial P^{\rho\sigma}}{\delta h^{\mu\nu}}D^{\mu\nu}_{\alpha}\eta^{\alpha}\bigg{)}\epsilon.$ (3.135)
@@
$\delta\widetilde{\eta}^{\alpha}-\widetilde{\delta}\widetilde{\eta}^{\alpha}=[\kappa^{2}Q^{\sigma}_{\tau}\partial_{\beta}\eta^{\tau}\eta^{\beta}-\kappa^{2}\partial_{\beta}(Q^{\sigma}_{\tau}\eta^{\tau})\eta^{\beta}-\kappa^{2}(\partial_{\tau}\eta^{\sigma})Q^{\tau}_{\beta}\eta^{\beta}-\kappa\frac{\delta Q^{\sigma}_{\tau}}{\delta h^{\mu\nu}}\eta^{\tau}D_{\beta}^{\mu\nu}\eta^{\beta}]\epsilon.$
@@
$+(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})[\widetilde{s}(\widetilde{h}^{\mu\nu})-s(\widetilde{h}^{\mu\nu})]-\kappa L_{\sigma}[\widetilde{s}\widetilde{\eta}^{\alpha}-s\widetilde{\eta}^{\alpha}]\bigg{]}.$ (3.136)
@@
Here $S^{rd}_{(0)gf}$ is the gauge fixed action but with the terms $sh^{\mu\nu}$ and $s\eta^{\alpha}$ replaced by $s\widetilde{h}^{\mu\nu}$ and $s\widetilde{\eta}^{\alpha}$, that is, with the ”non redefined” BRST variation of the ”redefined fields”. The important point is that the gauge fixed action
$S_{(0)gf}=\int d^{4}x[L_{s}+(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})D^{\mu\nu}_{\alpha}\eta_{0}^{\alpha}+\kappa^{2}L_{0\alpha}\partial_{\beta}\eta_{0}^{\alpha}\eta_{0}^{\beta}].$
@@
$S_{(0)gf}=\int d^{4}x[L_{s}+(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})sh^{\mu\nu}-L_{0\alpha}s\eta^{\alpha}].$
and which is clearly BRST invariant, may be replaced by a BRST invariant expression of the form
$S^{rd}_{(0)gf}=\int d^{4}x[L_{s}+(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})s(h^{\mu\nu}+P^{\mu\nu})-L_{0\alpha}s(\eta^{\alpha}-Q_{\beta}^{\alpha}\eta^{\beta})].$ (3.138)
This action is also BRST invariant since $s^{2}=0$ by an argument similar to (2.31) for Yang-Mills theories. It is in fact, not difficult to see that this argument applies to the present context. Alternatively, one may check this BRST invariance directly by calculating
@@
$\delta(\eta^{\alpha}-Q_{\beta}^{\alpha}\eta^{\beta})=\delta\eta^{\alpha}-\bigg{(}\frac{\partial Q^{\alpha}_{\beta}}{\partial h^{\mu\nu}}\delta h^{\mu\nu}\eta^{\beta}+Q_{\beta}^{\alpha}\delta\eta^{\beta}\bigg{)}$
@@
Having concluded that the action $S_{(0)gf}$ defined in (3.138) is BRST invariant, it is still not concluded that the functional $\Gamma_{1}$ defined in (3.137) is BRST invariant. What it can be said is that it is represented as the difference between two BRST invariant functionals. This imitates the Yang-Mills case, with the main difference that one functional is invariant under the BRST action $s$ while the other is invariant with respect to $\widetilde{s}$ defined in (3.134). The task is to elucidate if $\Gamma_{1}$ is approximately BRST invariant.
@@
$-\int d^{4}x[(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})sP^{\mu\nu}-L_{0\alpha}sQ^{\alpha}_{\beta}\eta^{\beta}]+O(\hbar^{2}).$ (3.140)
Note that this action contains only first order terms in $\hbar$. As for the Yang-Mills case, the previous action can be decomposed into a finite and divergent part by decomposing $P^{\mu\nu}=P_{f}^{\mu\nu}+P_{d}^{\mu\nu}$ and doing the same for $Q^{\alpha}_{\beta}$ and the parameters $\alpha$, $\beta$ and $\gamma$ of the gravitational part of the lagrangian. Following the procedure outlined in (2.88) for Yang-Mills fields, redefine
@@
$+\int d^{4}x[(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})sP_{d}^{\mu\nu}-L_{0\alpha}sQ^{\alpha}_{d\beta}\eta^{\beta}].$ (3.142)
@@
$+\int d^{4}x[(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})sP_{f}^{\mu\nu}-L_{0\alpha}sQ^{\alpha}_{f\beta}\eta^{\beta}].$
@@
$-\int d^{4}x[(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})sP_{f}^{\mu\nu}-L_{0\alpha}sQ^{\alpha}_{f\beta}\eta^{\beta}].$ (3.143)
@@
Furthermore, at the linear order in $\hbar$ one has that $P_{f}^{\mu\nu}(h^{\mu\nu})\sim P_{f}^{\mu\nu}(h^{\mu\nu}+P^{\mu\nu}_{f})$ as the difference is of second order. Also $Q_{f\beta}^{\alpha}(\eta^{\gamma})\sim Q_{f\beta}^{\alpha}(\eta^{\gamma}+Q^{\gamma}_{f\delta}\eta^{\delta})$. In addition $sP_{f}^{\mu\nu}\sim\widetilde{s}P_{f}^{\mu\nu}$ and $sQ^{\alpha}_{f\beta}\eta^{\beta}\sim\widetilde{s}Q^{\alpha}_{f\beta}\eta^{\beta}$ as the difference $\widetilde{s}-s$ is of first order in $\hbar$ and $P^{\mu\nu}$ and $Q^{\alpha}_{f\beta}\eta^{\beta}$ are also of that order, thus the difference affect only orders higher than linear. Therefore the last action can be written as

### v3

@@ 3.3 The propagator in certain higher order gravity theories
$S_{g}=\frac{1}{16\pi G_{N}}\int\bigg{[}R+RF_{1}(\square)R+R_{\mu\nu}F_{2}(\square)R^{\mu\nu}+R_{\mu\nu\alpha\beta}F_{3}(\square)R^{\mu\nu\alpha\beta}\bigg{]}\sqrt{-g}d^{4}x+S_{gf}.$ (3.107)
@@
$P^{2}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\theta_{\alpha\nu}+\theta_{\beta\nu}\theta_{\alpha\mu})-\frac{1}{3}\theta_{\beta\alpha}\theta_{\mu\nu},$
$P^{1}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\omega_{\alpha\nu}+\theta_{\beta\nu}\omega_{\alpha\mu}+\theta_{\alpha\mu}\omega_{\beta\nu}+\theta_{\alpha\nu}\omega_{\beta\mu}),$
@@
$=-k^{2}a(-k^{2})h^{\alpha\beta}(P^{2}_{\alpha\beta,\mu\nu}+P^{0}_{\alpha\beta,\mu\nu}+P^{1}_{\alpha\beta,\mu\nu}+P^{0-w}_{\alpha\beta,\mu\nu})h^{\mu\nu}.$
@@
$D^{\alpha\beta,\mu\nu}=-\frac{1}{k^{2}}\bigg{[}\frac{2}{a}P^{2}+2P^{1}+\frac{2a+d-1}{2D}P^{0-s}+\frac{a+3d}{2D}P^{0-w}+\frac{\sqrt{3}}{4D}(P^{0-sw}+P^{0-ws})\bigg{]}_{\alpha\beta,\mu\nu},$ (3.113)
@@ 3.4 Gravitational BRST symmetries
As for Yang-Mills fields, the gauge fixed action $S_{gf}$ is not invariant under the infinitesimal diffeomorphisms (3.101). In fact, it should not be so, as it reflects a particular choice of gauge. The task is to find a transformation, if it exists, leaving the gauge fixed action unchanged, perhaps up to a boundary term. This is the analog to the BRST symmetry described in (2.19)-(2.30). It is natural to try to generalize the procedure described in those formulas adapted to the present case. Instead of going through all the calculations as for the Yang-Mills case, one may employ the gained knowledge in those sections. A try may be to find a BRST transformation that increases the ghost number by one, supposing that these numbers are the same as for the Yang-Mills counterpart. In fact they are, as the assignation of the same ghost numbers than in Yang-Mills theories leads to a zero ghost number to the action $S_{gf}$, which is a desired feature. The ansatz for BRST transformation is then
$\delta h^{\mu\nu}=\kappa D_{\alpha}^{\mu\nu}\eta^{\alpha}\epsilon,\qquad\delta\eta^{\alpha}=-\kappa^{2}\partial_{\beta}\eta^{\alpha}\eta^{\beta}\epsilon,\qquad\delta\overline{\eta}_{\tau}=-\frac{\kappa^{3}}{\lambda}\square^{2}F_{\tau}\epsilon.$ (3.116)
@@
$\delta s\eta^{\alpha}=k^{4}\partial_{\beta}(\partial_{\gamma}\eta^{\alpha}\eta^{\gamma}\epsilon)\eta^{\beta}+\kappa^{4}\partial_{\beta}\eta^{\alpha}\partial_{\gamma}\eta^{\beta}\eta^{\gamma}\epsilon$
@@
$-\frac{\kappa^{3}}{\lambda}(\square^{2}F_{\tau}\epsilon)F_{\mu\nu}^{\tau}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\overline{\eta}_{\tau}F_{\mu\nu}^{\tau}\delta(D^{\mu\nu}_{\alpha}\eta^{\alpha}).$ (3.117)
@@
$\delta(D^{\mu\nu}_{\alpha}\eta^{\alpha})=-\kappa^{2}D^{\mu\nu}_{\alpha}\partial_{\gamma}\eta^{\alpha}\eta^{\gamma}\epsilon+\kappa\frac{\delta D_{\gamma}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\epsilon}^{\alpha\beta}\eta^{\epsilon}\eta^{\gamma}\epsilon$
$=-\kappa^{2}D^{\mu\nu}_{\alpha}\partial_{\gamma}\eta^{\alpha}\eta^{\gamma}\epsilon+\frac{\kappa}{2}\bigg{(}\frac{\delta D_{\epsilon}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\gamma}^{\alpha\beta}-\frac{\delta D_{\gamma}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\epsilon}^{\alpha\beta}\bigg{)}\eta^{\epsilon}\eta^{\gamma}\epsilon,$ (3.118)
@@
$(\delta_{\eta}\delta_{\xi}-\delta_{\xi}\delta_{\eta})h^{\mu\nu}=\bigg{(}\frac{\delta D_{\epsilon}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\gamma}^{\alpha\beta}-\frac{\delta D_{\gamma}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\epsilon}^{\alpha\beta}\bigg{)}\eta^{\epsilon}\xi^{\gamma}.$
@@
$\bigg{(}\frac{\delta D_{\epsilon}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\gamma}^{\alpha\beta}-\frac{\delta D_{\gamma}^{\mu\nu}}{\delta h^{\alpha\beta}}D_{\epsilon}^{\alpha\beta}\bigg{)}\eta^{\alpha}\eta^{\gamma}=2\kappa D^{\mu\nu}_{\alpha}\partial_{\gamma}\eta^{\alpha}\eta^{\gamma}.$
@@ 3.5 The effective gravitational action
In analogy with the procedure outlined in (2.33), the effective action $G(T,\xi,\overline{\xi},K,L)$ for gravitational theories, generating the connected correlation functions, is defined by the following relation
@@
$\exp\{i\int d^{4}x[L_{s}-\frac{\kappa^{2}}{2\lambda}F_{\tau}\square F^{\tau}+\overline{\eta}_{\tau}F_{\mu\nu}^{\tau}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa K_{\mu\nu}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa^{2}L_{\alpha}\partial_{\beta}\eta^{\alpha}\eta^{\beta}]\}.$ (3.119)
Here the current $J_{\mu\nu}h^{\mu\nu}$ is the analogous to the gauge field current coupling $J_{\mu}A^{\mu}$ for Yang-Mills theories, and $J_{\mu\nu}$ may be identified as the energy momentum tensor $J_{\mu\nu}=T_{\mu\nu}$ for an external source. As in those theories, the sources $K_{\mu\nu}$ and $L_{\alpha}$ are coupled to the nilpotent part of the BRST invariance. From this expression it follows that the gauge fixed action for the gravitational perturbation can be expressed as follows
@@ 3.6 Consequences of gravitational Slavnov-Taylor identiies
$\frac{\delta S_{(0)gf}}{\delta K_{\mu\nu}}=\kappa D^{\mu\nu}_{\alpha}\eta_{0}^{\alpha},\qquad\frac{\delta S_{(0)gf}}{\delta h^{\mu\nu}}=(\kappa K_{\alpha\beta}-\overline{\eta}^{\tau}\overline{F}_{\alpha\beta}^{\tau})\frac{\delta D_{\gamma}^{\alpha\beta}}{\delta h_{0}^{\mu\nu}}\eta_{0}^{\gamma}+\frac{\delta S_{0}}{\delta h_{0}^{\mu\nu}},$
@@
Here the reversed arrow $\overleftarrow{}$ indicates the action of the operator on the left. With the help of the last two formulas, the solution $\Gamma_{1}=S+\int d^{4}x\sigma_{0}X$, can be found by use of (3.131) and (3.129), the result is
@@ 3.7 Physical interpretation of the counter terms for gravitational theories
$+\int d^{4}x\bigg{[}-(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})\frac{\partial P^{\rho\sigma}}{\delta h^{\mu\nu}}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa L_{\sigma}\bigg{(}\frac{\delta Q^{\sigma}_{\tau}}{\delta h^{\mu\nu}}\eta^{\tau}D_{\beta}^{\mu\nu}-\kappa Q^{\sigma}_{\tau}\partial_{\beta}(\eta^{\tau})\bigg{)}\eta^{\beta}\bigg{]},$ (3.135)
@@
$\widetilde{\delta}(\widetilde{h}^{\mu\nu})-\delta(\widetilde{h}^{\mu\nu})=\widetilde{D}^{\mu\nu}_{\alpha}\widetilde{\eta}^{\alpha}\epsilon-D^{\mu\nu}_{\alpha}\eta^{\alpha}\epsilon-\frac{\partial P^{\mu\nu}}{\delta h^{\gamma\delta}}D^{\gamma\delta}_{\alpha}\eta^{\alpha}\epsilon$
$=\bigg{(}\frac{\delta D^{\mu\nu}_{\alpha}}{\delta h^{\rho\sigma}}\eta^{\alpha}P^{\rho\sigma}-D^{\mu\nu}_{\alpha}(Q^{\alpha}_{\tau}\eta^{\tau})-\frac{\partial P^{\mu\nu}}{\delta h^{\rho\sigma}}D^{\rho\sigma}_{\alpha}\eta^{\alpha}\bigg{)}\epsilon.$ (3.137)
@@
$\delta\widetilde{\eta}^{\alpha}-\widetilde{\delta}\widetilde{\eta}^{\alpha}=[\kappa^{2}Q^{\alpha}_{\tau}\partial_{\beta}\eta^{\tau}\eta^{\beta}-\kappa^{2}\partial_{\beta}(Q^{\alpha}_{\tau}\eta^{\tau})\eta^{\beta}-\kappa^{2}(\partial_{\tau}\eta^{\alpha})Q^{\tau}_{\beta}\eta^{\beta}-\kappa\frac{\delta Q^{\alpha}_{\tau}}{\delta h^{\mu\nu}}\eta^{\tau}D_{\beta}^{\mu\nu}\eta^{\beta}]\epsilon.$
@@
$+(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})[\widetilde{s}(\widetilde{h}^{\rho\sigma})-s(\widetilde{h}^{\rho\sigma})]-\kappa L_{\alpha}[\widetilde{s}\widetilde{\eta}^{\alpha}-s\widetilde{\eta}^{\alpha}]\bigg{]}.$ (3.138)
@@
Here $S^{rd}_{(0)gf}$ is the gauge fixed action but with the terms $sh^{\mu\nu}$ and $s\eta^{\alpha}$ replaced by $s\widetilde{h}^{\mu\nu}$ and $s\widetilde{\eta}^{\alpha}$, that is, with the ”non redefined” BRST variation of the ”redefined fields”. The important point is that not only the gauge fixed action
$S_{(0)gf}=\int d^{4}x[L_{s}+(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})D^{\rho\sigma}_{\alpha}\eta_{0}^{\alpha}+\kappa^{2}L_{0\alpha}\partial_{\beta}\eta_{0}^{\alpha}\eta_{0}^{\beta}],$
@@
$S_{(0)gf}=\int d^{4}x[L_{s}+(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})sh^{\rho\sigma}-L_{0\alpha}s\eta^{\alpha}].$
is BRST invariant, but any expression of the form
$S^{rd}_{(0)gf}=\int d^{4}x[L_{s}+(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})s(h^{\rho\sigma}+P^{\rho\sigma})-L_{0\alpha}s(\eta^{\alpha}-Q_{\beta}^{\alpha}\eta^{\beta})],$ (3.140)
has this invariant property as well. The BRST invariance of the last action follows from the fact that $s^{2}=0$, by an argument similar to (2.31) for Yang-Mills theories. It is in fact, not difficult to see that this argument applies to the present context. Alternatively, one may check this BRST invariance directly by calculating
@@
$\delta(\eta^{\alpha}-Q_{\beta}^{\alpha}\eta^{\beta})=\delta\eta^{\alpha}-\bigg{(}\frac{\partial Q^{\alpha}_{\beta}}{\partial h^{\mu\nu}}\delta h^{\mu\nu}\eta^{\beta}+Q_{\beta}^{\alpha}\delta\eta^{\beta}\bigg{)},$
@@
Having concluded that the action $S^{rd}_{(0)gf}$ defined in (3.140) is BRST invariant, it is still not concluded that the functional $\Gamma_{1}$ defined in (3.139) is BRST invariant. What it can be said is that it is represented as the difference between two BRST invariant functionals. This imitates the Yang-Mills case, however there is a big difference with that case since her one functional is invariant under the BRST action $s$ while the other is invariant with respect to $\widetilde{s}$ defined in (3.136). The task is to elucidate if $\Gamma_{1}$ is approximately BRST invariant, taking into account this property.
@@
$-\int d^{4}x[(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})sP^{\rho\sigma}-L_{0\alpha}sQ^{\alpha}_{\beta}\eta^{\beta}]+O(\hbar^{2}).$ (3.142)
Note that this action contains only first order terms in $\hbar$. As for the Yang-Mills case, the previous action can be separated into a finite and divergent part by decomposing
$P^{\mu\nu}=P_{f}^{\mu\nu}+P_{d}^{\mu\nu},$
and doing the same for $Q^{\alpha}_{\beta}$ and the parameters $\alpha$, $\beta$ and $\gamma$ of the gravitational part of the lagrangian. Following the procedure outlined in (2.88) for Yang-Mills fields, redefine
@@
$+\int d^{4}x[(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})sP_{d}^{\rho\sigma}-L_{0\alpha}sQ^{\alpha}_{d\beta}\eta^{\beta}].$ (3.144)
@@
$+\int d^{4}x[(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})sP_{f}^{\rho\sigma}-L_{0\alpha}sQ^{\alpha}_{f\beta}\eta^{\beta}].$
@@
$-\int d^{4}x[(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})sP_{f}^{\rho\sigma}-L_{0\alpha}sQ^{\alpha}_{f\beta}\eta^{\beta}].$ (3.145)
@@
Furthermore, at the linear order in $\hbar$ one has that $P_{f}^{\mu\nu}(h^{\mu\nu})\sim P_{f}^{\mu\nu}(h^{\mu\nu}+P^{\mu\nu}_{f})$ as the error is of second order. Also $Q_{f\beta}^{\alpha}(\eta^{\gamma})\sim Q_{f\beta}^{\alpha}(\eta^{\gamma}+Q^{\gamma}_{f\delta}\eta^{\delta})$. In addition $sP_{f}^{\mu\nu}\sim\widetilde{s}P_{f}^{\mu\nu}$ and $sQ^{\alpha}_{f\beta}\eta^{\beta}\sim\widetilde{s}Q^{\alpha}_{f\beta}\eta^{\beta}$ as the difference $\widetilde{s}-s$ is of first order in $\hbar$ and $P^{\mu\nu}$ and $Q^{\alpha}_{f\beta}\eta^{\beta}$ are also of that order, thus the difference affect only orders higher than linear. Therefore the last action can be written as